# Predicting the Geothermal Gradient in Colombia: a Machine Learning Approach

### v1

@@ 3.2 Model Building and Evaluation
The initial phase involved the partitioning of the dataset into training and testing subsets with an 80/20 split, a standardized step for model construction, and subsequent validation with independent test data. This split facilitates the proper evaluation of the model’s ability to generalize. Furthermore, we performed hyperparameter optimization to fine-tune the XGBoost model, enhancing its predictive accuracy.
Following the hyperparameter tuning, the XGBoost model was trained using a squared error loss function. The training process incorporated a five-fold grid search approach to fine-tune parameters such as shrinkage, maximum tree depth, and subsampling rate. This iterative refinement ensured the model’s robustness and adaptability.
The validation of the model used a quantitative framework influenced by Lösing & Ebbing (2021) and Rezvanbehbahani et al. (2017), focusing on absolute and relative error metrics. For the assessment of absolute error in the test set, we computed both the root mean square error (RMSE) and mean absolute error (MAE). Furthermore, normalized root mean square error (nRMSE) and normalized mean absolute error (nMAE) were calculated to provide standard and relative error measurements that are invariant to rescaling, facilitating comparative analysis.
@@ 4 Data and features
In this section, we explain the data used for constructing our predictive model. The target feature $y$ was the Apparent Geothermal Gradient (AGG) measurements from over 4000 well logs along the colombian territory, and for the predictor features $X$ we use certain geological and geophysical features which are linked to each AGG value.
@@ 4.1 Apparent Geothermal Gradient
Colombian AGG measurements were taken from the Alfaro et al. (2009) Colombian Geothermal Gradients map, which includes the data from over 4000 wells (Figure 0(a). The methodology they adopted for calculating the geothermal gradient for each well involved the following steps:
@@
    Temperature Measurement Correction: Various models and methods were used to correct the effects of drilling on formation temperatures. The most widely used method for correcting BHT data was the AAPG method, which compares BHT measurements with equilibrium temperatures in wells in Louisiana and West Texas. The correction was achieved using a third-order polynomial equation, with BHT correction as a function of depth.
@@
    Data Sources and Treatment: Data were obtained from the Finder and LogDB tables supplied by EPIS. The organisation of data involved forming pairs of depth-temperature data, and the selection process involved criteria for data quality. Units were converted from degrees Fahrenheit and feet to degrees Celsius and kilometres. Temperatures were corrected using the empirical method of AAPG.
@@
    Estimation of Individual Well Gradients: The maximum gradient within the upper crust was assumed to be vertical and calculated using the equation $G=\Delta T/\Delta P=(TBHT-Ts)/Z$, where $TBHT$ is the corrected bottom hole temperature, $Ts$ is the surface temperature, and $Z$ is the depth. Surface temperature was estimated based on correlation with elevation.
@@
To deal with underrepresented high-gradient values, we defined a high-gradient threshold of 33.86 °$\textdegree C/km$, which corresponds to the value at the 94th percentile (Figure 0(b)). We then calculated the ratio between the low (below the threshold) and high (above the threshold) gradient values and defined a weight of 8 for samples above the threshold of high gradient, which was adjusted based on model evaluation.
Furthermore, it is important to consider that the geothermal gradient data have an inherent bias: Although the measurements were made mainly in basins, a considerable amount of the data is close to Andean and volcanic areas, so measurements made in areas farther away tend to be underrepresented and affected the prediction for areas where there are no measurements. To correct for this bias, we defined each well located more than 50 km from the nearest volcano as ”nonvolcanic” and assigned a weigh of 4 to those wells.
Finally, we constrained the training and test sets to a specific range: from the 1st percentile value of 14.87 ${}^{\circ}$/km to the 99th percentile value of 43.80 ${}^{\circ}$/km (Figure 0(b)). This selection ensures that the model is trained and evaluated in a confidence range.
Figure 1: (a) Measured geothermal gradients in Colombia, (b) Histogram of apparent geothermal gradient values indicating a right-skewed distribution. Dashed lines indicated the representative percentiles used to train and test the model.
@@ 4.2 Geological and Geophysical Data
The AGG dataset includes latitude and longitude, for georeferencing well sites. This geospatial positioning is used to integrate the spatial variations of geothermal data into the model. Alongside these coordinates, topography and elevation play a significant role. The topographic elevation data, derived from well data and NASA’s global Earth Digital Elevation Model (DEM) as of 2021, provides a vertical reference point which is essential for geothermal gradient measuerements.
Furthermore, subsurface characteristics are another key aspect of the dataset. Insights into the thermal structure of the lithosphere and the behavior of heat at depth are garnered from the depths to the Moho discontinuity and the Curie isotherm. These are sourced from studies by Camacho et al. (2014) and Uieda & Barbosa (2017). Moho discontinuity and elevation data are strictly correlated to litosphere thickness, which is a great predictor for the goeothermal gradient in the continent (Kolawole & Evenick, 2023); Studies from Quintero et al (2019) and Ponce & Hernández (2014) show that there is an inverse correlation between the thermal gradient and the Curie isotherm.
Additionally, the presence of geophysical anomalies significantly influences the dataset. Magnetic anomalies, documented in the WDMAM V2 map by Dyment et al. (2020), are important as they potentially correlate with AGG. Deepest magnetic sources can be associated to Curie depth (Lösing & Ebbing, Op. Cit).
Gravity anomalies are strong indicators of the crustal composition as well. We use Free Air anomalies and Vertical Gravity Gradients from Sandwell et al. (2021) and Pavlis et al. (2012) datasets, it can indicate variations in the subsurface composition and structure. These variations are essential in understanding the distribution of subsurface heat and are influenced by crustal properties, such as the density variations and their correlation with rock types. The Bouguer anomaly was calculated using the Free Air anomaly with the Bouguer correction, and using this parameter allows us to reduce the elevation influence in the gravitational anomaly. After testing, we decided to keep the three gravity features.
@@
Lastly, the depth to basement rocks (igneous and metamorphic) is a critical factor in understanding the influence of subsurface rock types on thermal behavior, as they are a main source of radiogenic heat production compared to sedimentary rocks (Hasterok et al., 2018). Due to the absence of a specific basement depth model, an alternative method is employed using the geological map by Gómez et al. (2015). This method involves classifying rock outcrops as either basement or non-basement (sedimentary and deposits), and measuring the distance from each well site to the nearest basement outcrop. This assessment helps in understanding the impact of these rocks on geothermal gradients.
The integration of geospatial coordinates, topographical data, subsurface characteristics, geophysical anomalies, fault data, and the proximity to basement rocks forms a comprehensive framework for accurately modeling and understanding the subsurface geothermal behavior, reflecting the multidisciplinary nature of geothermal research. Each of these factors contributes uniquely to understand the gradient distribution, underlining the complex interplay of geological, geophysical, and spatial factors in geothermal exploration and modeling.
Detailed descriptions of these variables are provided in Table 1 and two graphical examples are shown in figure 2. The proper interpretation and utilization of this data are crucial for the predictive modeling of geothermal gradients across Colombia. The features were directly interpolated to obtain values at specific geographic points without resampling, the dataset with the lowest resolution (Moho Depth) has a 0.4° latitude-longitude grid.
@@
Elevation (m)
Height above sea level.
NASA Visible Earth (2021)
Moho Depth (m)
Depth to the Mohorovičić discontinuity.
Uieda & Barbosa (2017)
Curie Depth (Km)
Depth at which rocks in a specific geographical area encounter the Curie temperature
Camacho et al. (2014)
Magnetic Anomaly (nT)
Earth’s magnetic field anomaly at the point.
Dyment et al. (2020)
Free Air Anomaly (mGal)
Gravity anomaly after Free Air correction
Sandwell et al. (2021)
Bouguer Anomaly (mGal)
Gravity anomaly after Free Air correction and Bouguer corrections
Sandwell et al. (2021)
Vertical Gravity Gradient (E)
rate of change of vertical gravity ($gz$) with height ($z$)
Pavlis et al. (2012)
Faults (m)
Distance to the nearest normal, reverse or strike-slip fault.
Gómez et al., (2020)
Active Faults (m)
Distance to the nearest active fault.
Veloza et al., (2012)
Distance to basement (m)
Surface distance to igneous or metamorphic outcrops
Gómez et al., (2015)
Volcanic Domain
Boolean variable to define whether the well is located in a volcanic zone or not.
Gómez et al., (2020)
AGG measurements (°C/Km)
Target variable, measured from bottom hole data
Alfaro et al., (2009)
Figure 2: Side-by-side maps of Basement rocks outcrops and Bouguer anomaly with the measured gradient locations.
@@ 4.3 Hyperparameter tuning
To optimize the performance of the model, we defined a grid of hyperparameters. The code iterates through these hyperparameter combinations, training the model and evaluating mean squared error (MSE) on the test set for each configuration. The process systematically identifies the combination of hyperparameters that minimizes MSE, ensuring the model is fine-tuned for optimal predictive accuracy on the given dataset (Table 2).
@@
max_depth
7
Maximum depth of a tree.
learning_rate
0.05
Step size shrinkage used in updates.
gamma
0.15
Minimum loss reduction required to make a further partition.
min_child_weight
10
Minimum sum of instance weight needed in a child.
colsample_bytree
1
Fraction of features to be randomly sampled for each tree.
reg_alpha
0.5
L1 regularization term.
reg_lambda
2
L2 regularization term.
subsample
0.7
Fraction of training data to be randomly sampled for each boosting round.
@@ 5 Results
The result we present for this study is a predicted geothermal gradient map for all the Colombian territory. The accuracy of the methodology is evaluated using the test subset derived from the initial dataset partitioning. We examine the key performance metrics: MAE, RMSE, nMAE, nRMSE and the coefficient of determination ($R^{2}$) to quantify the predictive accuracy of the model. Additionally, we employ visualizations to facilitate a comprehensive analysis: plots for actual versus predicted values, diagrams illustrating feature importances, and assessments delineating the influence of individual features on the model’s output.
@@ 5.1 Model Validation
The initial phase of model evaluation involved scrutinizing its performance in both the training and the test subsets, the latter representing previously unseen data. For the training set, the model achieved an MAE of 1.8760, an RMSE of 2.5069, and a coefficient of determination ($R^{2}$) of 0.7670. These metrics are typical for models evaluated on their training data. Conversely, the test data got an MAE of 2.6130, an RMSE of 3.6430, and an $R^{2}$ of 0.5156.
@@
A detailed examination of the residuals, as depicted in Figure 3, reveals a random distribution around the zero line. This pattern suggests an absence of bias in the residuals, corroborated by a residual mean close to zero, indicating that the model is on average neither overestimating nor underestimating significantly. However, points with high residuals highlight instances where the model performance was suboptimal, though most residuals cluster near zero.
@@
Regarding the plot of actual versus predicted values, illustrated in Figure 4, the concentration of points along the ideal prediction line indicates proficient modeling of the target values. The dispersion of points as the target value increases suggests increasing variance in the model’s predictions. The model demonstrates higher accuracy for lower actual values while tending to underestimate higher actual values. The overlap of training (blue) and test (red) data points in the plot is indicative of good model generalization, evidenced by the similarity in performance across the training and test datasets. This overlap is a positive sign, suggesting that the model is not overfitting to the training data and most of the variance in the test set is towards the high gradient values.
@@
The map also shows regions where the model’s predictions deviate from measured data (as it’s evidenced in the Figure 4), those discrepancies suggest potential areas for model enhancement or may pinpoint inherent limitations within the data or the model’s structural assumptions. The tendency of the model to underestimate geothermal gradients, specially when they are higher, indicates possible gaps in its representation of heat source distribution or heat transfer mechanisms, and may require data enhancement.
@@ 5.2 Feature importance analysis
Feature importance provide insight into which variables are most informative for making accurate predictions. Multiple techniques were utilized to determine the importance of the features, each offering a unique perspective on the data.
@@
The weight-based feature importance, as shown in Figure 6, indicates that elevation and Moho depth are the most frequently used features in the model to make decisions; however, the importance is equilibrated. This highlights the model’s reliance on all the data to predict the target variable, especially topography, subsurface discontinuities, and structural influence.
@@
Figure 7 presents the gain-based feature importance, which reflects the improvement in accuracy brought by a feature to the branches, how the features contributed to minimize the loss function and, subsequently, to the model’s performance. Basement proximity in surface emerges as the most significant predictor, however, most of the features importances are equilibrated.
@@
Subsequently, the model performance was assessed as a function of the number of features used (Starting with the less important according to Feature Importance), as depicted in Figure 10. The impact of each feature on the training set (model score) and the test set ($R^{2}$ score) increases with the number of features, indicating improved precision and variance explanation, respectively. However, these scores flatten after including approximately nine features, suggesting that additional features would not contribute significantly to the predictive capability of the model regarding error metrics. However, it is important to recognize that each feature plays a crucial role in enhancing the model’s spatial accuracy, and omitting any feature could therefore significantly affect the outcomes.
@@
While some features have significant predictive power, the model reaches a performance threshold beyond which additional features do not yield further benefits. This information is important for refining the model to focus on the most impactful features and avoid overfitting by excluding less informative ones.

### v7

@@ 4.2 Model Building and Evaluation
The initial phase involved partitioning the data set into training and testing subsets with an 80/20 split, a standardized step for model construction, and subsequent validation with independent test data. This split facilitates the proper evaluation of the model’s ability to generalize. Furthermore, we performed hyperparameter optimization to fine-tune the XGBoost model, enhancing its predictive accuracy.
Following the hyperparameter tuning, the XGBoost model was trained using a squared error loss function. The training process incorporated a five-fold grid search approach to fine-tune parameters such as shrinkage, maximum tree depth, and subsampling rate. This iterative refinement ensured the robustness and adaptability of the model.
The validation of the model used a quantitative framework influenced by Lösing & Ebbing (2021) and Rezvanbehbahani et al. (2017), focusing on absolute and relative error metrics. For the assessment of the absolute error in the test set, we computed the mean square error (RMSE) and mean absolute error (MAE). Furthermore, normalized root mean square error (nRMSE) and normalized mean absolute error (nMAE) were calculated to provide standard and relative error measurements that are invariant to rescaling, facilitating comparative analysis.
@@ 5 Data and features
In this section, we explain the data used to build our predictive model. The target feature $y$ was the Apparent Geothermal Gradient (AGG) measurements from over 4000 well logs along the colombian territory, and for the predictor features $X$ we use certain geological and geophysical features which are linked to each AGG value.
@@ 5.1 Apparent Geothermal Gradient
Colombian AGG measurements were taken from the Alfaro et al. (2009) Colombian Geothermal Gradients map, which includes the data from over 4000 wells (Figure 5(a)). The methodology they adopted for calculating the geothermal gradient for each well involved the following steps:
@@
    Temperature Measurement Correction: Various models and methods were used to correct for the effects of drilling on formation temperatures. The most widely used method for correcting BHT data was the AAPG method, which compares BHT measurements with equilibrium temperatures in wells in Louisiana and West Texas. The correction was achieved using a third-order polynomial equation, with BHT correction as a function of depth.
@@
    Data Sources and Treatment: Data were obtained from the Finder and LogDB tables supplied by EPIS. The organization of data involved forming pairs of depth-temperature data, and the selection process involved criteria for data quality. Units were converted from degrees Fahrenheit and feet to degrees Celsius and kilometers. Temperatures were corrected using the empirical method of AAPG.
@@
    Estimation of Individual Well Gradients: The maximum gradient within the upper crust was assumed to be vertical and calculated using the equation $G=\Delta T/\Delta P=(TBHT-Ts)/Z$, where $TBHT$ is the corrected bottom hole temperature, $Ts$ is the surface temperature, and $Z$ is the depth. The surface temperature was estimated on the basis of correlation with elevation.
@@
Note that the gradient values obtained by Alfaro et al (2009) represent the average gradient over certain depth measured for each well. These apparent gradients are important for indicating broader regional trends and identifying areas with geothermal anomalies (Figure 4a). However, they may not precisely reflect localized thermal behaviors at specific depths. This is an important consideration for interpreting gradient data and applying it in our predictive model.
@@
Figure 5: (a) Measured geothermal gradients in Colombia, (b) Histogram of apparent geothermal gradient values indicating a right-skewed distribution. Dashed lines indicated the representative percentiles for lowest and extreme gradients that will be discarded.
@@
To deal with underrepresented high-gradient values, we defined a frequency-based weighting so that the less-frequent gradients within a 15-bin histogram distribution have the higher weights, and the most-frequent gradients have the lowest weight (Figure 6). The number of bins and, therefore, the values of the weights; were adjusted based on model evaluation.
Figure 6: 15-bin frequency histogram showing the weighting distribution for the geothermal gradients. The weights were normalized so the highest weight is 1.0 and the lowest is 0.02.
Furthermore, it is important to consider that the geothermal gradient data have an inherent bias: Although the measurements were made mainly in basins, a considerable amount of the data is close to Andean and volcanic areas, so measurements made in areas farther away tend to be underrepresented and affected the prediction for areas where there are no measurements. To correct for this bias, we defined each high-gradient well located more than 75 km from the nearest volcano as ”non-volcanic” and assigned a weigh of 8 to those wells.
Finally, we constrained the training and test sets to a specific range: from the 1st percentile value of 14.87 C/km to the 99th percentile value of 43.80 C/km (Figure 5(b)). This selection ensures that the model is trained and evaluated in a confidence range.
@@ 5.2 Geological and Geophysical Data
The AGG dataset includes latitude and longitude, for georeferencing well sites. This geospatial positioning is used to integrate the spatial variations of the geothermal data into the model. Alongside these coordinates, topography and elevation play a significant role. The topographic elevation data, derived from well data and NASA’s global Earth Digital Elevation Model (DEM) as of 2021, provide a vertical reference point which is essential for geothermal gradient measurements.
Additionally, subsurface characteristics are another key aspect of the dataset. Insights into the thermal structure of the lithosphere and the behavior of heat at depth are garnered from the depths to the Moho discontinuity and the Curie isotherm. These are sourced from studies by Quintero et al. (2014) and Uieda & Barbosa (2017). Moho discontinuity and elevation data are strictly correlated to litosphere thickness, which is a great predictor for the goeothermal gradient in the continent (Kolawole & Evenick, 2023); Studies from Quintero et al (2019) and Ponce & Hernández (2014) show that there is an inverse correlation between the thermal gradient and the Curie isotherm. The depth of lithosphere-asthenosphere boundary was regarded as one of the top five important features in previous models (e.g., Rezvanbehbahani et al., 2017; Lösing and Ebbing, 2021), however, it was discarded because of the lack of a representative dataset for Colombia.
Additionally, the presence of geophysical anomalies significantly influences the dataset. Magnetic anomalies, documented in the WDMAM V2 map by Dyment et al. (2020), are important as they potentially correlate with AGG. The deepest magnetic sources can be associated to the Curie depth (Lösing & Ebbing, 2021).
Gravity anomalies are also strong indicators of crustal composition. We use Free Air anomalies and Vertical Gravity Gradients from Sandwell et al. (2021) and Pavlis et al. (2012) datasets, it can indicate variations in the subsurface composition and structure. These variations are essential in understanding the distribution of subsurface heat and are influenced by crustal properties, such as the density variations and their correlation with rock types. The Bouguer anomaly was calculated using the Free Air anomaly with the Bouguer correction, and using this parameter allows us to reduce the elevation influence in the gravitational anomaly. After testing, we decided to keep the three gravity features.
@@
Lastly, the depth to the basement rocks (igneous and metamorphic rocks) is a critical factor in understanding the influence of subsurface rock types on thermal behavior, as they are a main source of radiogenic heat production compared to sedimentary rocks (Hasterok et al., 2018). Due to the absence of a specific basement depth model, an alternative method is employed using the geological map by Gómez et al. (2015). This method involves classifying rock outcrops as either basement (igneous and metamorphic) or non-basement (sedimentary and deposits), and measuring the distance from each well site to the nearest basement outcrop. This assessment, along with other features such as faulting or structural constraints that may affect the heat flow, helps to understand the impact of these rocks on geothermal gradients.
Integration of geospatial coordinates, topographical data, subsurface characteristics, geophysical anomalies, fault data, and proximity to basement rocks forms a comprehensive framework for accurately modeling and understanding the subsurface geothermal behavior, reflecting the multidisciplinary nature of geothermal research. Each of these factors contributes uniquely to understand the gradient distribution, underlining the complex interplay of geological, geophysical, and spatial factors in geothermal exploration and modeling.
Detailed descriptions of these variables are provided in Table 1 and two graphical examples are shown in Figure 7. The correct interpretation and use of these data is crucial for predictive modeling of geothermal gradients across Colombia. The features were directly interpolated to obtain values at specific geographic points without resampling, the dataset with the lowest resolution (Moho Depth) has a 0.4° latitude-longitude grid.
@@
Elevation (m) Height above sea level. NASA Visible Earth (2021)
Moho Depth (m) Depth to the Mohorovičić discontinuity. Uieda & Barbosa (2017)
Curie Depth (Km) Depth at which rocks in a specific geographical area encounter the Curie temperature Quintero et al. (2014)
Magnetic Anomaly (nT) Earth’s magnetic field anomaly at the point. Dyment et al. (2020)
Free Air Anomaly (mGal) Gravity anomaly after Free Air correction Sandwell et al. (2021)
Bouguer Anomaly (mGal) Gravity anomaly after Free Air correction and Bouguer corrections Sandwell et al. (2021)
Vertical Gravity Gradient (E) rate of change of vertical gravity ($gz$) with height ($z$) Pavlis et al. (2012)
Faults (m) Distance to the nearest normal, reverse or strike-slip fault. Gómez et al., (2020)
Active Faults (m) Distance to the nearest active fault. Veloza et al., (2012)
Distance to basement (m) Surface distance to igneous or metamorphic outcrops Gómez et al., (2015)
Volcanic Domain Boolean variable to define whether the well is located in a volcanic zone or not. Gómez et al., (2020)
AGG measurements (°C/Km) Target variable, measured from bottom hole data Alfaro et al., (2009)
Figure 7: Side-by-side maps of Basement rocks outcrops and Moho Depth with the measured gradient locations.
@@ 5.3 Hyperparameter tuning
To optimize the performance of the model, we defined a grid of hyperparameters. The code iterates through these hyperparameter combinations, training the model, and evaluating mean squared error (MSE) on the test set for each configuration. The process systematically identifies the combination of hyperparameters that minimizes MSE, ensuring the model is fine-tuned for optimal predictive accuracy on the given dataset (Table 2).
@@
max_depth 10 Maximum depth of a tree.
learning_rate 0.1 Step size shrinkage used in updates.
gamma 0.0 Minimum loss reduction required to make a further partition.
min_child_weight 1 Minimum sum of instance weight needed in a child.
colsample_bytree 0.5 Fraction of features to be randomly sampled for each tree.
reg_alpha 0.5 L1 regularization term.
reg_lambda 0.5 L2 regularization term.
subsample 1 Fraction of training data to be randomly sampled for each boosting round.
@@ 6 Results
The result we present for this study is a predicted geothermal gradient map for all the Colombian territory. The accuracy of the methodology is evaluated using the test subset derived from the initial dataset partitioning. We examine the key performance metrics: MAE, RMSE, nMAE, nRMSE and the coefficient of determination ($R^{2}$) to quantify the predictive precision of the model. Additionally, we employ visualizations to facilitate a comprehensive analysis: plots for actual versus predicted values, diagrams illustrating feature importances, and assessments delineating the influence of individual features on the model’s output.
@@ 6.1 Model Validation
The initial phase of the model evaluation involved scrutinizing its performance in both the training and the test subsets, the latter representing previously unseen data. For the training set, the model achieved an MAE of 1.4793, an RMSE of 2.1371, and a coefficient of determination ($R^{2}$) of 0.8287. These metrics are typical for models evaluated on their training data. Conversely, the test data got an MAE of 2.6826, an RMSE of 3.5842, and an $R^{2}$ of 0.5497.
@@
A detailed examination of the residuals, as depicted in Figure 8, reveals a random distribution around the zero line. This pattern suggests an absence of bias in the residuals, corroborated by a residual mean close to zero, indicating that the model is on average neither overestimating nor underestimating significantly. However, points with high residuals highlight instances where model performance was suboptimal, though most residuals cluster near zero.
@@
Regarding the plot of actual versus predicted values, illustrated in Figure 9, the concentration of points along the ideal prediction line indicates proficient modeling of the target values. The dispersion of points as the target value increases suggests increasing variance in the model’s predictions. The model demonstrates higher accuracy for lower actual values while tending to underestimate higher actual values in the test set. The overlap of training (blue) and test (red) data points in the plot is indicative of good model generalization, evidenced by the similarity in performance across the training and test datasets. This overlap is a positive sign, suggesting that the model is not overfitting to the training data and most of the variance in the test set is towards the high gradient values.
@@
The Figure 11 highlights the differences between the actual and the predicted geothermal gradient values, showing regions where the model’s predictions deviate from measured data. Those discrepancies suggest potential areas for model enhancement or may pinpoint inherent limitations within the data or the model’s structural assumptions. The tendency of the model to underestimate geothermal gradients, specially when they are higher, indicates possible gaps in its representation of heat source distribution or heat transfer mechanisms, and may require data enhancement.
Figure 11: Differences between actual and model-predicted geothermal gradient values.
@@ 6.2 Feature importance analysis
Feature importance provide insight into which variables are most informative for making accurate predictions. Multiple techniques were used to determine the importance of the features, each offering a unique perspective on the data.
@@
The weight-based feature importance, as shown in Figure 12, indicates that active faults, elevation and Moho depth are the most frequently used features in the model to make decisions, with structural features and basement proximity having also a high importance. Potential fields and Curie Depth, on the other hand, are less frequently used. The importance is equilibrated among most of the features. This highlights the model’s reliance on all the data to predict the target variable, especially topography, subsurface discontinuities, and structural influence.
@@
Figure 13 presents the gain-based feature importance, which reflects the improvement in accuracy brought by a feature to the branches, how the features contributed to minimize the loss function, and, subsequently, to the model’s performance. Basement proximity in surface emerges as the most significant predictor, and features such as Potential fields and Curie Depth play a significant role here as well; however, most of the features importances are equilibrated.
@@
Assessing the model’s performance as a function of the feature count, starting with the less important (gain-based) and so on, reveals an increase in both the model score for the training set and the $R^{2}$ score for the test set, as shown in Figure 16. These increments signify enhanced precision and variance explanation. Nonetheless, the progressive improvement in scores flatten after incorporating roughly nine features, indicating that further features may not substantially enhance the model’s error metrics. It is essential to note, however, that each feature significantly enhances the model’s spatial precision, as suggested by the feature importance and mutual information data. Thus, excluding any feature could negatively affect the outcomes: while performance levels off after nine features, the remaining features should not necessarily be dismissed. Incorporation of particularly insightful features, like the actual depth to the basement, depth of lithosphere-asthenosphere boundary or improved feature resolution could, indeed, offer advancements to the model.